1 pt a) Find the magnitude of -A + B, where A-(35.1, 67.3), B = (89.9,-156.0). 8. AO 60.274BO 75.342 DO 117.722 EO 147.153 FO 183.941 GO 229.926 HO 287.407 CO 94.178 pbFind the direction of -A + B, where A - (35.1, 67.3), B- (89.9,-156.0). Express the angle counterclockwise relative to the positive r-axis. in deg) 9. AO 145.300BO 181.625CO 227.031 DO 283.788 EO 354.736FO 443.419 GO 554.274 HO 692.843
Answer #1

Solution,

8)

-A = -35.1i - 67.3j

B = 89.9i - 156j

-A + B = 54.8i - 223.3j

Magnitude = sqrt(54.8^2 + 223.3^2) = 229.926 (G)

9)

Direction, theta = arctan (-223.3/54.8) = - 76.21

Direction = 360 - 76.21 = 283.788 (D)

Comment in case any doubt please rate my answer ....